OpenStudy (anonymous):

please help, fan and medal (: Henry rolls 2 number cubes numbered 1 through 6 while playing his favorite board game. He will get a second turn if he rolls a sum that is an even number less than 10. What are Henry's chances of getting a second turn when he rolls the number cubes? seven over eighteen 11 over 18 5 over 36 17 over 36

OpenStudy (anonymous):

number cubes?

OpenStudy (anonymous):

I think it's 17/36 but hang on a second so I can see if I can explain it

OpenStudy (anonymous):

on planet earth they are called "dice" and your job is to count the number of ways you can roll a number less than 10, i.e. the number of ways you can roll 2, 4, 6, or 8

OpenStudy (anonymous):

do you have a dice table? that would help immensely

OpenStudy (anonymous):

Yes add up all the numbers, and then select the ones out of the total numbers (select the evens under 10) and that should give you 17/36

OpenStudy (anonymous):

i can see if i can find one for you the number of ways is as follows: 1 way to get a 2 3 ways to get a 4 5 ways to get a 6 5 ways to get an 8

OpenStudy (anonymous):

It hsould only take you about 5 minutes to add up all the numbers and count. You don't really need a dice table

OpenStudy (anonymous):

which in my calculation to not add up to 17 ways, but rather to 14 ways

OpenStudy (anonymous):

But that wasn't an option. Are you sure @satellite73

OpenStudy (anonymous):

I'm so confused x.x

OpenStudy (anonymous):

go with \[\frac{14}{36}=\frac{7}{18}\] hold on let me find you a table

OpenStudy (anonymous):

Oh I see

OpenStudy (anonymous):

sit tight give me a second

OpenStudy (anonymous):

your right I found the table.

OpenStudy (anonymous):

that one you have to fill in for yourself but it would be good practice and should not take longer than 3 minutes

OpenStudy (anonymous):

OMG . YOURE FREAKING AMAZING

OpenStudy (anonymous):

Nice job satellite! You explained it to me too

OpenStudy (anonymous):

here is one that is filled in if you want to do nothing

OpenStudy (anonymous):

but you will remember better if you fill it in yourself the number of ways to roll a 7 is 6, so the probability you roll a 7 is \(\frac{6}{36}=\frac{1}{6}\) that is the most likely roll fill in the table and keep it, then you can answer any probability question about dice by referring to the table and counting

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

yw